Consider the second order partial differential equation du/dt= d^2u/dx^2 +2du/dx+u over the domain x in [0,l) and t>=0. It is given that u(0,t)=u(l,t)=0. Use the method of separation of variables to prove that the general solution with the given boundary condition is u(x,t)= infinity series n=1 bnsin(npix/l)exp(-x-((npi/l)^2)t) where bn is a constant for every n N  

Hint u(x,t)=X(x)T(t)tnsit te Seind ond partial difertinl cuatan +2n St the dowain e To,e) an Use metod o Separet ion Vaiades to rore tht the He g
Answer #1

Seconol adeor pasti al oiffen ent Oa Given Sodution is Eauation au +u +2 het the So Qution ulxt = X (1) T(4) ul,t = XT -Q alAgaim Take 2 dx X -K d2x d2 dx Auxiliay aon The is m 2m K = o -2 ±J4 -4k 2 JK-I -i + l So e C Ca JK- x S Jk-i x put get eoeJK-I=hT JK-1 2 2 TT 2 T2 put 4 Sho h77 So The genenal Soluton s b S 11 nsl ohene Amner
Answer #2
d^2u/dt^2=c^2d^2u/dx^2. 00 Boundary conditions. u(0.t)=u(l.t)=0 Initial conditions: u (x.0)=sin(πx/l) du/dt(x.0)=0 Solve the equation using the separation of variables method.
answered by: Rayene